# Fast Inference of Removal-Based Node Influence

### 13 Mar 2024 v1

@@ Abstract.
Graph neural networks (GNNs) are widely utilized to capture the information spreading patterns in graphs. While remarkable performance has been achieved, there is a new trending topic of evaluating node influence. We propose a new method of evaluating node influence, which measures the prediction change of a trained GNN model caused by removing a node. A real-world application is, “In the task of predicting Twitter accounts’ polarity, had a particular account been removed, how would others’ polarity change?”. We use the GNN as a surrogate model whose prediction could simulate the change of nodes or edges caused by node removal. To obtain the influence for every node, a straightforward way is to alternately remove every node and apply the trained GNN on the modified graph. It is reliable but time-consuming, so we need an efficient method. The related lines of work, such as graph adversarial attack and counterfactual explanation, cannot directly satisfy our needs, since they do not focus on the global influence score for every node. We propose an efficient and intuitive method, NOde-Removal-based fAst GNN inference (NORA), which uses the gradient to approximate the node-removal influence. It only costs one forward propagation and one backpropagation to approximate the influence score for all nodes. Extensive experiments on six datasets and six GNN models verify the effectiveness of NORA. Our code is available at https://github.com/weikai-li/NORA.git.
@@ 1. Introduction
Figure 1. An example of the task-specific influence of node removal in social networks. Red versus Blue represents two different opinions, and color shades represent the degree of opinion. When the top blue node is removed, the two pink nodes hear less voice from the blue nodes and become red. The two left nodes no longer follow the middle node, and the left white node becomes blue. These are the influence of removing the top blue node.
They define the node influence based on the graph structure (e.g., connectivity) or a given propagation model (e.g., susceptible-infected-removed model). However, these definitions are not flexible enough to capture the node influence from different aspects. For example, we might want to identify the biggest political influencers on Twitter. In this case, we want to calculate the influence score of a Twitter account based on how much it would affect other users’ political polarity had it been removed. In another scenario, we might want to identify the biggest fashion influencers on Twitter, and we want to calculate the node influence based on how much it would affect other users’ fashion categories had it been removed. Instead of adopting any fixed node influence definition, we thus focus on task-specific node influence score calculation based on node removal. Figure 1 provides an example of the task-specific node influence.
@@
  • Methodology. We propose an efficient and effective algorithm, NORA, to approximate the node influence for all nodes.
@@ 2.2. Graph Counterfactual Explanation
Graph counterfactual explanation aims to explain why a GNN model gives a particular result of a target node/edge/graph. Unlike the factual explanation that explains by associating the prediction with critical nodes or edges, the counterfactual explanation tries to find the minimum perturbation on the input graph that can change the prediction of the target. Some methods (Ying et al., 2019; Lucic et al., 2021; Bajaj et al., 2021; Luo et al., 2020) are purely based on edge removal; some methods utilize both node removal and edge removal, and the methods include optimizing mask matrices (Tan et al., 2022; Wu et al., 2021), predicting node influence (Rong et al., 2023), applying graph generation models (Ma et al., 2022; Yuan et al., 2020), or searching for an optimal neighbor graph (Yuan et al., 2021; Huang et al., 2023). As analyzed in Section 1 and shown in Table 1, our problem setting differs from existing works, so these methods are not directly applicable to our problem setting. We modify and adapt two widely used methods as supplementary baselines to this new problem. The first baseline is inspired by optimizing a mask matrix, which is a common method in graph counterfactual explanation (Lucic et al., 2021; Ying et al., 2019; Tan et al., 2022; Bajaj et al., 2021; Wu et al., 2021). They usually multiply the adjacency matrix with the mask matrix which indicates edge existence. During training, its elements are within $[0,1]$. During inference, elements below $0.5$ indicate edge removal. We adapt it to node removal by using a node mask vector. Our second baseline is inspired by a recent work, LARA (Rong et al., 2023), which trains a GCN model to predict node influence on the explanation target. The parameter size does not grow with the input graph size, so it is more scalable compared to previous methods. We adapt it as our second baseline.
@@ 3.2. Problem Definition
(3) $\mathrm{F}_{g_{\theta}}(v_{r})=\sum_{e_{ij}\in D_{e}-D_{r}}||f_{e}(g_{\theta}(G)_{e_{ij}})-f_{e}(g_{\theta}(G_{-v_{r}})_{e_{ij}})||_{1}\,,$
@@ 4. Methods
The task is to predict the defined influence score for every node on the graph. The biggest challenge is efficiency, and we want to avoid traversing all the nodes. We propose an intuitive, simple, yet effective method, NOde-Removal-based fAst GNN inference (NORA). In general, we approximate the node influence by analyzing the calculation formula and decomposing it into three parts, which correspond to three kinds of influence of the node removal. Then, we use gradient information and some heuristics to approximate it. Figure 3 illustrates the three kinds of influence. We mainly introduce our method in the node classification task, and after that, we will explain how to generalize it to the link prediction task.
@@ 4.2. Approximation of Each Decomposed Term
(11) $T_{2}\approx k_{3}\cdot\delta Topo_{r},\delta Topo_{r}=\sum_{i\in\mathcal{N}(r)}\sum_{j\in\mathcal{N}(i)}\hat{\delta}\alpha_{ji},$
@@
$T_{3}$: Change of hidden representations of other nodes. When we analyze the term $T_{3}$, on the computation graph, $\boldsymbol{h_{j}^{(L-1)}}$ can reach $f_{r}$ by either $\boldsymbol{h_{r}^{(L)}}$ or $\boldsymbol{h_{i}^{(L)}},i\neq r$. Based on the chain rule, we can simplify $T_{3}$ in Equation 7 to Equation 12, which can be further transformed into Equation 13.
@@
(14) $T_{3}\approx\big{(}\sum_{i\neq r}\boldsymbol{\frac{\partial f_{r}}{\partial h_{i}^{(L-1)}}}\boldsymbol{\delta h_{i}^{(L-1)}}\big{)}\big{(}1-\frac{d_{r}}{(N-1)(d+\beta)}\big{)}.$
@@ 5.1. Baseline Adaption
Table 4. Pearson correlation coefficient between real influence scores and approximated scores on citation datasets.
@@
Table 5. Pearson correlation coefficient between real influence scores and approximated scores on Twitter datasets.
@@ 5.5. Time Consumption
Table 6. Time of calculating the ground truth.
@@ Appendix B Supplementary Experiments
B.1. Enlarging Label Usage of Baseline
@@
Table 8. Performance with more labels.
@@ B.4. Experiment Details
We also tune the hyper-parameters for NORA and every baseline method, and we tune them for each dataset and each GNN model. NORA has five hyper-parameters: $k_{1}$, $k_{2}$, $k_{2}^{\prime}$, $k_{3}^{\prime}$, and $\beta$. Among them, $\beta$ is related to approximating term $T_{1}$ and term $T_{3}$; $k_{1}$, $k_{2}$, and $k_{2}^{\prime}$ are related to approximating term $T_{2}$; $k_{3}^{\prime}$ is used to adjust the weight between the two components in Equation 16. For convenience, we normalize the two components before multiplying $k_{3}^{\prime}$. Based on our experience, we can separately tune the hyper-parameters for the two components, and the best hyper-parameter setting is usually the combination of the best hyper-parameter setting for the two components. $\beta$ is usually within [1, 20]; $k_{1}$, $k_{2}$, and $k_{2}^{\prime}$ are within [0, 1]; $k_{3}^{\prime}$ is usually within [0.5, 5] and not far from one, since we normalize the two components.

### 16 Mar 2024 v2

@@ Abstract.
Graph neural networks (GNNs) are widely utilized to capture the information spreading patterns in graphs. While remarkable performance has been achieved, there is a new trending topic of evaluating node influence. We propose a new method of evaluating node influence, which measures the prediction change of a trained GNN model caused by removing a node. A real-world application is, “In the task of predicting Twitter accounts’ polarity, had a particular account been removed, how would others’ polarity change?”. We use the GNN as a surrogate model whose prediction could simulate the change of nodes or edges caused by node removal. Our target is to obtain the influence score for every node, and a straightforward way is to alternately remove every node and apply the trained GNN on the modified graph to generate new predictions. It is reliable but time-consuming, so we need an efficient method. The related lines of work, such as graph adversarial attack and counterfactual explanation, cannot directly satisfy our needs, since their problem settings are different. We propose an efficient, intuitive, and effective method, NOde-Removal-based fAst GNN inference (NORA), which uses the gradient information to approximate the node-removal influence. It only costs one forward propagation and one backpropagation to approximate the influence score for all nodes. Extensive experiments on six datasets and six GNN models verify the effectiveness of NORA. Our code is available at https://github.com/weikai-li/NORA.git.
@@ 1. Introduction
Figure 1. An example of the task-specific influence of node removal in social networks. Red versus Blue represents two different opinions, and color shades represent the degree of opinion. When the top blue node is removed, the two pink nodes might hear less voice from the blue nodes and become red. The two left nodes might no longer follow the middle node, and the left white node might become blue. These are the influence of removing the top blue node.
They define the node influence based on the graph structure (e.g., connectivity) or a designed information propagation model (e.g., the susceptible-infected-removed model). However, these definitions are not flexible enough to capture the node influence from different aspects. For example, we might want to identify the biggest political influencers on Twitter. In this case, we want to calculate the influence score of a Twitter account based on how much it would affect other users’ political polarity had it been removed. In another scenario, we might want to identify the biggest fashion influencers on Twitter, and we want to calculate the node influence based on how much it would affect other users’ fashion categories had it been removed. Instead of adopting any fixed node influence definition, we thus focus on task-specific node influence score calculation based on node removal. Figure 1 provides an example of the task-specific node influence.
@@
  • Methodology. We propose an efficient, intuitive, and effective algorithm, NORA, to approximate the node influence for all nodes.
@@ 2.2. Graph Counterfactual Explanation
Graph counterfactual explanation aims to explain why a GNN model gives a particular result of a target node/edge/graph. Unlike the factual explanation that explains by associating the prediction with critical nodes or edges, the counterfactual explanation tries to find the minimum perturbation on the input graph that can change the prediction of the target. Some methods (Ying et al., 2019; Lucic et al., 2021; Bajaj et al., 2021; Luo et al., 2020) are purely based on edge removal; some methods utilize both node removal and edge removal, and the methods include optimizing mask matrices (Tan et al., 2022; Wu et al., 2021), predicting node influence (Rong et al., 2023), applying graph generation models (Ma et al., 2022; Yuan et al., 2020), or searching for an optimal neighbor graph (Yuan et al., 2021; Huang et al., 2023). As analyzed in Section 1 and shown in Table 1, our problem setting differs from existing works, so these methods are not directly applicable. We modify and adapt two widely used graph counterfactual explanation methods as the baseline methods. The first baseline is inspired by optimizing a mask matrix (Lucic et al., 2021; Ying et al., 2019; Tan et al., 2022; Bajaj et al., 2021; Wu et al., 2021). The mask matrix indicates edge existence, and it is multiplied by the adjacency matrix before the message passing. Its elements are within $[0,1]$, and the mask matrix is optimized to maximize the difference between original predictions and new predictions. We adapt it to node removal by using the node mask vector. Our second baseline is inspired by LARA (Rong et al., 2023), which trains a GCN model to predict node influence on the explanation target. The parameter size does not grow with the input graph size, so it is more scalable compared to previous methods.
@@ 3.2. Problem Definition
(3) $\mathrm{F}_{g_{\theta}}(v_{r})=\sum_{e_{ij}\in D_{e}-D_{r}}|f_{e}(g_{\theta}(G)_{e_{ij}})-f_{e}(g_{\theta}(G_{-v_{r}})_{e_{ij}})|\,,$
@@ 4. Methods
Our target is to calculate the defined influence score for every node on a graph. The biggest challenge is efficiency, and we want to avoid traversing all the nodes. We propose an intuitive yet effective method, NOde-Removal-based fAst GNN inference (NORA). In general, we approximate the node influence by analyzing the calculation formula and decomposing it into three parts, which correspond to three kinds of influence of the node removal. Then, we use gradient information and some heuristics to approximate it. Figure 3 illustrates the three kinds of influence. We mainly introduce our method in the node classification task, and after that, we will explain how to generalize it to the link prediction task.
@@ 4.2. Approximation of Each Decomposed Term
(11) $T_{2}\approx k_{3}\cdot\delta Topo_{r},\quad\delta Topo_{r}=\sum_{i\in\mathcal{N}(r)}\sum_{j\in\mathcal{N}(i)}\hat{\delta}\alpha_{ji},$
@@
$T_{3}$: Change of hidden representations of other nodes. When we analyze the term $T_{3}$, on the computation graph, $\boldsymbol{h_{j}^{(L-1)}}$ can reach $\boldsymbol{f_{r}}$ by either $\boldsymbol{h_{r}^{(L)}}$ or $\boldsymbol{h_{i}^{(L)}},i\neq r$. Based on the chain rule, we can simplify $T_{3}$ in Equation 7 to Equation 12, which can be further transformed into Equation 13.
@@
(14) $T_{3}\approx T_{3}^{\prime}=\big{(}\sum_{i\neq r}\boldsymbol{\frac{\partial f_{r}}{\partial h_{i}^{(L-1)}}}\boldsymbol{\delta h_{i}^{(L-1)}}\big{)}\big{(}1-\frac{d_{r}}{(N-1)(d+\beta)}\big{)}.$
@@ 5.1. Baseline Adaption
Table 4. Pearson correlation coefficient between real influence and approximated influence on citation datasets.
@@
Table 5. Pearson correlation coefficient between real influence and approximated influence on Twitter datasets.
@@ 5.5. Time Consumption
Table 6. Time of calculating the real influence.
@@ Appendix B Supplementary Experiments
B.1. Enlarging Label Usage of the “Prediction” Method
@@
Table 8. Performance of the “prediction” method with more label usage.
@@ B.4. Experiment Details
We also tune the hyper-parameters for NORA and every baseline method, and we tune them for each dataset and each GNN model. NORA has five hyper-parameters: $k_{1}$, $k_{2}$, $k_{2}^{\prime}$, $k_{3}^{\prime}$, and $\beta$. Among them, $\beta$ is related to approximating term $T_{1}$ and term $T_{3}$; $k_{1}$, $k_{2}$, and $k_{2}^{\prime}$ are related to approximating term $T_{2}$; $k_{3}^{\prime}$ is used to adjust the weight between the two components in Equation 16. For convenience, we normalize the two components before multiplying $k_{3}^{\prime}$. Based on our experience, we can separately tune the hyper-parameters for the two components, and the best hyper-parameter setting is usually the combination of the best hyper-parameter setting for the two components. $\beta$ is usually within [1, 20]; $k_{1}$, $k_{2}$, and $k_{2}^{\prime}$ are within [0, 1]; $k_{3}^{\prime}$ is usually within [0.5, 5] and not far from 1, since we normalize the two components.